OpenStudy (anonymous):

URGENT I NEED HELP ON MY ASSIGNMENT! 6 QUESTIONS! WILL FAN AND MEDAL!

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14 Question 3 (Multiple Choice Worth 2 points) (03.07) Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x - 7 y = - one halfx - 2 y = 2x - 2 y = - one halfx - 7 Question 4 (Multiple Choice Worth 1 points) (03.07) Graph with a line going through points negative 2 comma zero and 0 comma negative one. Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1 Question 5 (Multiple Choice Worth 2 points) (03.07) Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = -one halfx + 3 y = -one halfx + 6 y = 2x + 3 Question 6 (Multiple Choice Worth 1 points) (03.07) Graph with a line going through points zero comma two point five and four comma two point five. Select the equation of the line that passes through the point (3, -1) and is parallel to the line on the graph. y = -1 y = 3 y = x -1 y = 3x - 1 Question 7 (Multiple Choice Worth 1 points) (03.07) Select the coordinates of two points that are on the line x = 5. (5, 0) and (0, 5) (5, 0) and (5, -2) (5, 0) and (0, 2) (0, 5) and (0, 2)

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

FLVS